At January 1, 2020, Buffalo Company’s outstanding shares included the following.

298,000 shares of $50 par value, 7% cumulative preferred stock
921,000 shares of $1 par value common stock


Net income for 2020 was $2,560,000. No cash dividends were declared or paid during 2020. On February 15, 2021, however, all preferred dividends in arrears were paid, together with a 5% stock dividend on common shares. There were no dividends in arrears prior to 2020.

On April 1, 2020, 426,000 shares of common stock were sold for $10 per share, and on October 1, 2020, 106,000 shares of common stock were purchased for $21 per share and held as treasury stock.

Compute earnings per share for 2020. Assume that financial statements for 2020 were issued in March 2021. (Round answer to 2 decimal places, e.g. $2.55.)

Earnings per share

$
Answer #1

The earnings per share = Net income available for common shareholders/ Weighted average number of shares

= $1,517,000/1,274,700 = 1.19

The earnings per share = $1.19

Notes

1) Net income available for common shareholders = Net Income - Preferred Share Dividend

= $2,560,000 - (298,000 shares*$50)*7%

= $2,560,000 -$1,043,000

Net income available for common shareholders = $1,517,000

2) Weighted Average of shares as of Mar 31 2021

Beginning balance = (921,000)/4 = 230,250

Shares issued = (921,000+426,000)/2 = 673,500

Shares bought back = (921,000+426,000-106,000)/4 = 310,250

Weighted average no. of shares = 230,250+673,500+310,250 = 1,214,000

Stock dividend on common shares =5%

Weighted average no. of shares = 1,214,000*(1+5%)

Weighted average no. of shares = 1,274,700